A social scientist claims that only 70% of all high school seniors capable of doing college work actually go to college. If a sample of 15 people are chosen, compute the mean and the standard


deviation.


1
Expert's answer
2021-05-07T12:38:56-0400

We have given that,

p = 70% = 0.7

q = 0.3

n = 15

Using the binomial dstribution,

Mean "= np = 15 \times 0.7 = 10.5"

Standard deviation "= \sqrt{npq} = \sqrt{15\times 0.7\times 0.3} = 1.77"